OpenStudy (anonymous):

can someone help me with solving linear equations with substitution...please?!

OpenStudy (lalaly):

sure

OpenStudy (anonymous):

cool thanks...\[3x-2y=-3 \] \[x+5y=-18\]

OpenStudy (amistre64):

what do you believe it means to "substitute"?

OpenStudy (anonymous):

do I have to rewrite one of them?

OpenStudy (lalaly):

amistre can help :D:D

OpenStudy (anonymous):

I have to fine the value for y or x and substitute it in the equation

OpenStudy (amistre64):

not quite a value, more like a suitable "value" so that either the y or x speaks the same language ...

OpenStudy (anonymous):

ok so looking at the x+5y=-18 if I "flop" the x to get the 5y to = the rest of the equation...is that the right idea?

OpenStudy (amistre64):

you usually try to pick one of the equations and solve for a variable that looks easiest to do ... and if i read you right, yes x + 5y = -18 ; lets solve for "x" x = -5y -18 is a suitable value to plug into the other equation

OpenStudy (anonymous):

so it can be wither the x or y?

OpenStudy (amistre64):

3x -2y = -3 ; but we know that x has to equal some value that is akin to -5y-18 so lets substitute that "value" and solve for y

OpenStudy (amistre64):

it can be whatever variable you are most comfortable with solving for

OpenStudy (anonymous):

and now I plug it into the the first equation for x...?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

3x - 2y = -3 3(-5y-18) -2y = -3 , we have "substituted" and can solve for "y" now

OpenStudy (anonymous):

ok Im working it...ty!

OpenStudy (amistre64):

youre welcome ; if you get stuck, just let us know :)

OpenStudy (anonymous):

y=3

OpenStudy (amistre64):

it might .... lets see 3(-5y-18) -2y = -3 -15y -46 -2y = -3 -15y -2y = 43 y(-15-2) = 43 y = 43/-17 ... did i do it right?

OpenStudy (amistre64):

i tend to perfer elimination over substitution; less work involved: 3x −2y=−3 x +5y=−18 ; *-3 3x −2y=−3 -3x-15y =54 ------------ y(-17) = 51 y = -51/17 2 17 3 ---- 51 y = -3 is it

OpenStudy (amistre64):

i had forgotten how to multiply 3 and -17 lol

OpenStudy (anonymous):

I got y=-3, great! Thanks again! I just wasn't sure which one to flop I thought it was always the y! I forget signs all the time...

OpenStudy (amistre64):

..... oy vey, your gonna have to just correct my typos

OpenStudy (anonymous):

and with elimination I would of just found a number that would of canceled something out between the two...correct?

OpenStudy (amistre64):

yep; elimination is "eliminating" a variable by modifying the equations thru a suitable scalar multiple

OpenStudy (anonymous):

and if the problem said use 'addition"?

OpenStudy (amistre64):

thats the same as elimination; they like to throw different terms at you that amount to the same thing

OpenStudy (amistre64):

by adding the equations toghter you eliminate one of the variables after youve scaled one of them

OpenStudy (anonymous):

such as \[2x-3y=14\] and\[4x-6y=28\] I could use -2 on the top one?

OpenStudy (amistre64):

-2 would work great for the top; or even -1/2 on the bottom

OpenStudy (anonymous):

yeah im not fond of fractions...lol

OpenStudy (amistre64):

if you simply multiply the top by 2 you will see that the equations are the same

OpenStudy (anonymous):

so addition does mean elimination...? I appreciate your help!

OpenStudy (amistre64):

addition is another name for elimination yes

OpenStudy (anonymous):

ok cool got it. thanks again!